OpenStudy (anonymous):

(x)/(x^2+9x+20) - 4/(x^2+7x+12) subtracting rational expressions

OpenStudy (anonymous):

idk why but i get x+20 / (x+3) (x+4) (x+5)

OpenStudy (lgbasallote):

\[\implies \frac{x}{(x+5)(x+4)} - \frac{4}{(x+4)(x+3)}\] so LCD is (x+5)(x+4)(x+3) \[\implies \frac{x(x+3) - 4(x+5)}{(x+5)(x+4)(x+3)}\] \[\implies \frac{x^2 + 3x - 4x - 20}{(x+5)(x+4)(x+3)}\] \[\implies \frac{x^2 - x - 20}{(x+5)(x+4)(x+3)}\] \[\implies \frac{(x-5)(x+4)}{(x+5)(x+4)(x+3)}\] does that give you an idea what to next?

OpenStudy (anonymous):

how did u get the (x-5) (x+4) on the numerator on the last step?

OpenStudy (anonymous):

wouldnt it be (x-20) since u combine like terms and subtract x from x^2 which gave us x-20

OpenStudy (lgbasallote):

you cant subtract x from x^2

OpenStudy (lgbasallote):

x^2 and x are different terms

OpenStudy (anonymous):

oh ok

OpenStudy (lgbasallote):

so x^2 - x -20 stays that way

OpenStudy (lgbasallote):

then factor it out becomes (x-5)(x+4)

OpenStudy (anonymous):

ohhhhhhhhh

OpenStudy (anonymous):

thanks

OpenStudy (lgbasallote):

welcome ^_^

OpenStudy (lgbasallote):

well i need to go now so i hope you can proceed on your own :D

OpenStudy (anonymous):

yes thank u so much!!

OpenStudy (lgbasallote):

wonderful